Peak expiratory (PEF) is a measure of a patient's ability to excel air from the lungs. Patients with asthma or other respiratory conditions often have restricted PEF. An investigator conducts a study to investigate whether there is a difference in mean PEF in children with chronic bronchitis as compared to those without. Data on PEF are collected and summarized below. Based on the data, is there statistical evidence of a lower mean PEF in children with chronic bronchitis as compared to those without? Run the appropriate test at significance level =0.05

Group A chronic bronchits # of children= 42 Mean PEF= 265 Std Dev PEF= 65

Group B No chronic bronchitis # of children 40 Mean PEF 319 Std Dev PEF 70

a)Hypthoses HO: vs HA b) Caluculate S2p= c) Compute test stat d) P-value e) Conclusion Accept HO or Reject HO f) Calculate a 95% confidence interval for mean difference
Answer #1

(a)

H0: Null Hypothesis:

HA: Alternative Hypothesis:

(b)

(c)

So,

Test statistic is:

t = (265- 319)/14.9091 = - 3.62

(d)

ndf = n1 + n2 - 2 = 42 + 40 - 2 = 80

By Technology, P - Value = 0.000257

(e)

Since P - value = 0.000257 is less than = 0.05, the difference is significant. Reject null hypothesis.

Correct option:

Reject H0

(f)

From Table, critical values of t = 1.99

Confidence Interval:
(265 - 319) (1.99 X 14.9091)

=- 54 29.6691

= (- 83.6691, - 24.3309)

So,

Confidence Interval:

- 83.6691 < < - 24.3309